Commissioning Notes for 20080424 UT

Summary:

Running Notes/Log (UTtimes):

Open at sunset. Problems with air for SX M1, MCSPU

03:35 Everything up and running...

WT10_399 Winds from S, pointing to NE. Pointing off 8.0 -15.0 (thermal) ~0.8" FWHM

03:55 AGW homed, ready to test GCS. Because of telescope thermal equalibriam pointing problems the on axis was off the image plane. Probe movement was done by hand to put the probe on axis. Guide images acquired and sub imaged. Images numbered 372 to 436. Guiding stopped. Guide image off 2 arcsec from wfs. Image was driven out of field of view in x direction.

Sign change was implemented.

04:23 Attempted guiding again and successfully closed the loop!

04:35 Set PA to 90. Still moving stage by hand as pointing is off by 14 arcsecs. Move Stage. Attempting to close to loop. Loop closed. At PA of 90 we spiraled away from the star. Reason for spiral is under investigation.

04:51 While waiting for spiral solution will attempt to determine North on the guider CCD. Set PA to 0. For PA of 0 determined West is to the left.

05:04 Perform same experiment for PA of 90. For PA of 90 determined West is up.

05:09 Going back to PA of 0. Noticed oscillation. Attempted to collect data and before successful the oscillations stopped. Dave T thinks the oscillations are greater than 3 Hz because the star appears to change size (focus) of one arcsec in .1 second.

05:30 Attempt to guide on axis. PA set to 0. Preset MODE = GUIDE. It sent telescope, moved acquisition camera on axis. Closed the loop. Guiding!!!

05:37 Stopped guiding by issueing another preset with PA = 90. Will attempt to guide again.

05:37 Attempt to guide with PA set to 90 on axis. Spiraled off of star (expected behavior) as we have not yet implemented PA not 0 fix.

05:42 Attempt off axis guiding to test if GCS can properly move the guide probe. Send preset of PA = 314.09, predicted position for stage should be 7.5 in x, and 453 in y. Guider did not move to correct coordinates. Previous probe position was 7.485, 600.50. Following this attempt it is 7.182, 600.503. Probe is still on axis. Investigation is underway.

06:46 Another attempt at off axis guiding but with PA = 0. The coordinates it sent the guider to were on axis. This is not correct. Was delivered a huge guide offset that caused the telescope to slew. Guide offsets should be capped at 10 arcsecs in order to avoid bogus guide offsets.

06:55 It appears slewing was caused by inverted scale factor using 1/correct-scale factor. PCS is being modified to provide correct scale factor. Scale factor is 1/focal length, where focal length is 123769.1 mm at F/15.

07:02 New PCS started.

07:03 Try PA=.01, off axis guide star test again. Did not find star. Determined sign problem on x axis is wrong because of tertiary mirror. This problem will not show up at Direct Gregorian. Were to make sign correction is being investigated.

Other point of interest. While performing a manual positioning in x, y of the probe, Dave received a problem with AGw UMAC. Problem statement was "call failed while checking bits on UMAC: RPC: timed out." This may have happened because the GCS and manual usage of the "setxy -x 21.97 -y 502.94" may have been issuing commands simultaneously to the AGw. Torsten will test this on Thursday during the day to determine if he can cause similar behavior. This was written up as IssueTrak # 1297

07:40 Updated GCS to fix multiple errors (plate scale, sign application, ...).

07:43 Try test again with updated GCS. Guiding star ran out of field of view. Problem being investigated.

07:55 Updated from absolute coordinates to deltas in GCS.

08:00 Try test again with updated GCS. Guiding star ran out of field of view. Problem being investigated. Investigation shows the GCS is performing the correct guide offsets to PCS. PCS is driving away from star such that star leaves field of view. PCS is being investigated.

Starting Thursday, Torsten will be working in the late day too early evening on some GCS enhancements for robustness and some additional cmd line commands to make the commissioners job easier.

08:30 Noticed vibrating IRTC images.

We have taken three series of images while problem was occurring. These are located in /Repository or in /Repository/20080424_IRTC with timestamps of 083007 cube 00007 (Az = 30, El = 78, windspeed 7.6 m/s from Az 193, temp 2.3 C, mirror ventilation was on, AGw was on), turned mirror ventilation off without improvement, turned ventilation back on, stoped AGw without improvement, turned AGw back on, took image sequences 083804 cube 00008, and 084031 cube 00009.

mpeg movies of fits files. DLM 25Apr08

08:45 Before attempting to turn-off treehouse cooling fans we moved to zenith and back to the target to see if the problem continues or goes away. The problem continued. Will slow down the four treehouse fans. Disabling mount. Treehouse fans put on slow speed.

08:55 Re-enabling mount. Will go back to target to see if problem continues or is gone. With the treehouse fans turned down it appears the vibration is less (perhaps gone). NOTE: Before turning fans on slow vibration was very obvious. With fans on slow vibration is not obvious. Image sequences captured are 085730 cube 00010, 085856 cube 00011, 090351 cube 00012 (seeing was very bad so it may not be very usable), 090718 cube 00013 (Az = -6, El = 80, wind 9.4 m/s from Az = 200, temp = 2.2 C so conditions where mostly consistent since beginning of sequence), 090934 cube 00014.

09:26 Observing different star. Seeing has worsened. Star does not appear to be vibrating. Star is faint.

09:27 Taking images for EF pair.

OSS was not responding. amp fault, bad angles. Had to kill OSS and restart. Hexapod was than responsive.

*******************************************************************************************************************

11:00 The 4 am polynomial drop out occurred. Michele was running top and Norm tried to grab the processes running at the time:

[ncushing@lbtmu102 ~]$ ps -ef
UID PID PPID C STIME TTY TIME CMD
root 1 0 0 2007 ? 00:00:13 init [5]
root 2 1 0 2007 ? 00:00:02 [migration/0]
root 3 1 0 2007 ? 00:00:04 [ksoftirqd/0]
root 4 1 0 2007 ? 00:00:02 [migration/1]
root 5 1 0 2007 ? 00:00:04 [ksoftirqd/1]
root 6 1 0 2007 ? 00:00:00 [events/0]
root 7 1 0 2007 ? 00:00:00 [events/1]
root 8 6 0 2007 ? 00:00:00 [khelper]
root 9 6 0 2007 ? 00:00:00 [kacpid]
root 38 6 0 2007 ? 00:00:00 [kblockd/0]
root 39 6 0 2007 ? 00:00:00 [kblockd/1]
root 40 1 0 2007 ? 00:00:00 [khubd]
root 52 6 0 2007 ? 00:00:00 [aio/0]
root 53 6 0 2007 ? 00:00:00 [aio/1]
root 51 1 0 2007 ? 00:06:03 [kswapd0]
root 126 1 0 2007 ? 00:00:00 [kseriod]
root 202 1 0 2007 ? 00:00:00 [scsi_eh_0]
root 203 1 0 2007 ? 00:00:00 [ahc_dv_0]
root 212 1 0 2007 ? 00:00:00 [scsi_eh_1]
root 213 1 0 2007 ? 00:00:00 [ahc_dv_1]
root 228 6 0 2007 ? 00:00:00 [kmirrord]
root 229 6 0 2007 ? 00:00:00 [kmir_mon]
root 241 1 0 2007 ? 02:51:43 [kjournald]
root 1208 1 0 2007 ? 00:00:00 udevd
root 1795 6 0 2007 ? 00:00:00 [kauditd]
root 1918 1 0 2007 ? 00:00:00 [kjournald]
root 2534 1 0 2007 ? 03:05:53 syslogd -m 0
root 2538 1 0 2007 ? 00:00:00 klogd -x
root 2549 1 0 2007 ? 00:00:30 irqbalance
rpc 2561 1 0 2007 ? 00:00:00 portmap
rpcuser 2581 1 0 2007 ? 00:00:00 rpc.statd
root 2616 1 0 2007 ? 00:00:00 rpc.idmapd
root 2649 1 0 2007 ? 00:00:33 [rpciod]
root 2650 1 0 2007 ? 00:00:00 [lockd]
root 2778 1 0 2007 ? 00:02:01 /usr/sbin/automount --timeout=60 /misc file /etc/auto.misc
root 2824 1 0 2007 ? 00:00:49 /usr/sbin/automount --timeout=60 /home ldap lbtmu111.lbt.as.arizona.edu:ou=auto.home,dc=lbto,dc=org
nscd 2883 1 0 2007 ? 00:00:20 /usr/sbin/nscd
root 2900 1 0 2007 ? 00:00:00 /usr/sbin/smartd
root 2910 1 0 2007 ? 00:00:00 /usr/sbin/acpid
root 2964 1 0 2007 ? 00:00:13 /usr/sbin/sshd
root 2979 1 0 2007 ? 00:00:00 xinetd -stayalive -pidfile /var/run/xinetd.pid
ntp 2992 1 0 2007 ? 00:00:10 ntpd -u ntp:ntp -p /var/run/ntpd.pid -g
root 3011 1 0 2007 ? 00:00:02 sendmail: accepting connections
smmsp 3019 1 0 2007 ? 00:00:00 sendmail: Queue runner@01:00:00 for /var/spool/clientmqueue
root 3030 1 0 2007 ? 00:00:00 gpm -m /dev/input/mice -t imps2
root 3040 1 0 2007 ? 00:00:00 crond
xfs 3072 1 0 2007 ? 00:00:00 xfs -droppriv -daemon
root 3091 1 0 2007 ? 00:00:00 /usr/sbin/atd
dbus 3110 1 0 2007 ? 00:00:00 dbus-daemon-1 --system
root 3120 1 0 2007 ? 00:00:00 cups-config-daemon
root 3131 1 0 2007 ? 00:41:49 hald
root 3327 1 0 2007 tty1 00:00:00 /sbin/mingetty tty1
root 3330 1 0 2007 tty2 00:00:00 /sbin/mingetty tty2
root 3338 1 0 2007 tty3 00:00:00 /sbin/mingetty tty3
root 3346 1 0 2007 tty4 00:00:00 /sbin/mingetty tty4
root 3348 1 0 2007 tty5 00:00:00 /sbin/mingetty tty5
root 3350 1 0 2007 tty6 00:00:00 /sbin/mingetty tty6
root 3351 1 0 2007 ttyS0 00:00:00 /sbin/agetty -L 9600 ttyS0 vt102
root 3352 1 0 2007 ttyS1 00:00:00 /sbin/agetty -L 9600 ttyS1 vt102
root 3354 1 0 2007 ? 00:00:00 /usr/bin/gdm-binary -nodaemon
root 3779 3354 0 2007 ? 00:00:00 /usr/bin/gdm-binary -nodaemon
root 3782 3779 0 2007 ? 00:13:48 /usr/X11R6/bin/X :0 -audit 0 -auth /var/gdm/:0.Xauth -nolisten tcp vt7
gdm 3852 3779 0 2007 ? 00:02:25 /usr/bin/gdmgreeter
root 32051 7 0 Apr15 ? 00:00:01 [pdflush]
root 32052 6 0 Apr15 ? 00:00:03 [pdflush]
root 17680 2964 0 Apr16 ? 00:00:00 sshd: root@pts/3
root 17685 17680 0 Apr16 pts/3 00:00:00 -bash
root 6631 1 0 Apr20 ? 00:00:03 cupsd
root 6668 1 0 Apr20 ? 00:00:00 /usr/sbin/snmpd -Lsd -Lf /dev/null -p /var/run/snmpd -a
root 9133 2964 0 Apr22 ? 00:00:00 sshd: delapena [priv]
delapena 9136 9133 0 Apr22 ? 00:00:00 sshd: delapena@pts/6
delapena 9137 9136 0 Apr22 pts/6 00:00:00 -tcsh
1033 5339 1 0 Apr23 ? 00:00:00 networkserver
1033 5343 1 0 Apr23 ? 00:00:01 rpcserver ok
1033 5345 1 1 Apr23 ? 00:06:52 gshmserver ok
root 8697 2964 0 Apr23 ? 00:00:00 sshd: delapena [priv]
delapena 8703 8697 0 Apr23 ? 00:00:00 sshd: delapena@pts/4
delapena 8705 8703 0 Apr23 pts/4 00:00:00 -tcsh
1033 10019 1 0 00:02 ? 00:01:35 PCS
root 12549 3040 0 04:02 ? 00:00:00 crond
root 12550 12549 0 04:02 ? 00:00:00 /bin/bash /usr/bin/run-parts /etc/cron.daily
root 12551 12550 0 04:02 ? 00:00:00 /usr/bin/perl -w /etc/cron.daily/00-logwatch
root 12552 12550 0 04:02 ? 00:00:00 awk -v progname=/etc/cron.daily/00-logwatch progname {????? print progname ":\n"????? progname="";???? }???? { print;
root 12783 12551 0 04:07 ? 00:00:00 /bin/mail -s LogWatch for lbtmu102.lbt.as.arizona.edu root
delapena 12794 9137 0 04:07 pts/6 00:00:00 top
root 12828 2964 0 04:08 ? 00:00:00 sshd: ncushing [priv]
ncushing 12831 12828 0 04:08 ? 00:00:00 sshd: ncushing@pts/2
ncushing 12832 12831 0 04:08 pts/2 00:00:00 -tcsh
root 12865 12551 0 04:08 ? 00:00:00 sh -c ( /bin/cat /tmp/logwatch.QBX12556/messages | /etc/log.d/scripts/shared/onlyservice 'shaperd\d' |/etc/log.d/scripts/shared/remoroot 12866 12865 0 04:08 ? 00:00:00 sh -c ( /bin/cat /tmp/logwatch.QBX12556/messages | /etc/log.d/scripts/shared/onlyservice 'shaperd\d' |/etc/log.d/scripts/shared/remoroot 12867 12866 2 04:08 ? 00:00:00 /bin/cat /tmp/logwatch.QBX12556/messages
root 12868 12866 87 04:08 ? 00:00:02 /usr/bin/perl -w /etc/log.d/scripts/shared/onlyservice shaperd\d
root 12869 12866 0 04:08 ? 00:00:00 /usr/bin/perl -w /etc/log.d/scripts/shared/removeheaders
root 12870 12866 0 04:08 ? 00:00:00 /usr/bin/perl -w /etc/log.d/scripts/services/shaperd
ncushing 12872 12832 1 04:08 pts/2 00:00:00 ps -ef

*******************************************************************************************************************

Polynomials eventually received and able to slew to target.

11:13 PCS issues discussed with David Terrett during the past couple of hours. As a result the PCS was updated.

11:15 Attempting to guide off axis. MCSPU is not seeing polynomials. Problem is persistent and being investigated.

11:26 PCS updated. Issue test preset from Preset GUI. Problem still occurring. Stopping PCS and MCSPU.

11:30 Restart MCSPU. Restart PCS. Send test preset from IRTC. No polynomials

11:42 Decided to call it a night and regroup late in the afternoon on the 24th to readdress the PCS polynomial situation.

-- DavidThompson & NormCushing - 24 Apr 2008

I Attachment Action Size Date Who Comment
irtc_00007_cube.mpegmpeg irtc_00007_cube.mpeg manage 1 MB 24 Apr 2008 - 20:34 DougMiller movie: irtc_00007_cube.mpeg
irtc_00008_cube.mpegmpeg irtc_00008_cube.mpeg manage 1 MB 24 Apr 2008 - 20:40 DougMiller movie: irtc_00008_cube.mpeg
irtc_00009_cube.mpegmpeg irtc_00009_cube.mpeg manage 2 MB 24 Apr 2008 - 20:41 DougMiller movie: irtc_00009_cube.mpeg
irtc_00010_cube.mpegmpeg irtc_00010_cube.mpeg manage 1 MB 24 Apr 2008 - 20:41 DougMiller movie: irtc_00010_cube.mpeg
irtc_00011_cube.mpegmpeg irtc_00011_cube.mpeg manage 1 MB 24 Apr 2008 - 20:41 DougMiller movie: irtc_00011_cube.mpeg
irtc_00012_cube.mpegmpeg irtc_00012_cube.mpeg manage 2 MB 24 Apr 2008 - 20:50 DougMiller  
irtc_00013_cube.mpegmpeg irtc_00013_cube.mpeg manage 1 MB 25 Apr 2008 - 02:32 DougMiller movie: irtc_00013_cube.mpeg
irtc_00014_cube.mpegmpeg irtc_00014_cube.mpeg manage 1 MB 25 Apr 2008 - 04:09 DougMiller movie: irtc_00014_cube.mpeg
Topic revision: r8 - 25 Apr 2008, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback